OpenStudy (peachtea):

Which ordered pairs are on the line with equation ​ y+6.75=0.25(x−1)? Select each correct answer. ​ (2, −6) ​ (7, −5) ​ (4, −6)​ ​ (−3, −7.75)

OpenStudy (justin234):

I believe A and C

OpenStudy (peachtea):

Are you sure? :/

OpenStudy (justin234):

if u didnt do it already yes lol :3

OpenStudy (peachtea):

Okay okay, thaaaank you, I hope it's right :)

OpenStudy (peachtea):

Oh really?

OpenStudy (justin234):

really what?

OpenStudy (lonelyy__):

yes. it would be (2, 6.5) on the graph

OpenStudy (peachtea):

Oh okay thank you :o

OpenStudy (justin234):

what no medal for me

OpenStudy (smartperson):

no

OpenStudy (lonelyy__):

b is correct

OpenStudy (justin234):

lame

OpenStudy (peachtea):

sorry :(

OpenStudy (smartperson):

i already gave the medal though

OpenStudy (lonelyy__):

and so is (:

OpenStudy (justin234):

well unmedal her and medal me were both right and i was first

OpenStudy (justin234):

she has 2 and i got none

OpenStudy (lonelyy__):

and so is the last one

OpenStudy (lonelyy__):

so your answer would be b,c, and d :))

OpenStudy (lonelyy__):

thank you for the medal

OpenStudy (lonelyy__):

there you go @Justin234

OpenStudy (justin234):

thanks and there u go

OpenStudy (lonelyy__):

thank you

OpenStudy (justin234):

no prob

OpenStudy (shadowlegendx):

For the individual who looks this question up in a search engine and wants to see it solved out... \[y+6.75=0.25(x−1)\] This is point slope form, let simplify it to slope intercept form. \[y + 6.75 = \frac{ 1 }{ 4 }x - 0.25\] Subtract 6.75 from both sides \[y = \frac{ 1 }{ 4 }x - 7\]

OpenStudy (shadowlegendx):

So we are at (0, -7) on the graph. Following the slope in either direction Positive: (4, -6) Negative: (-4, -8) Positive means it rose by \[\frac{ 1 }{ 4 }\] Negative means it rose by \[\frac{ -1 }{ -4 }\]

OpenStudy (shadowlegendx):

So the answer is C

OpenStudy (peachtea):

Only C?

OpenStudy (peachtea):

@ShadowLegendX

OpenStudy (justin234):

so I was right C and A

OpenStudy (shadowlegendx):

Can't be ​ (−3, −7.75) as it's not starting at a decimal, but (0, -7) It's also not rising or running by decimals. It also can't be ​ (2, −6) as if the run is 4 ( where the rise is 1, from 1/4 and the run is 4) Because it would go 4 units in either direction in relation to the x axis, not 2 units.

OpenStudy (shadowlegendx):

Actually, it isn't (2, −6) ​ (7, −5) It is ​ (4, −6)​ and probably ​ (−3, −7.75) as well. Since it rises 3/4 of the way, or goes down -0.75 of a unit Then runs 3/4 of the way, or by 3 units to the left

OpenStudy (peachtea):

Thank you :)

OpenStudy (shadowlegendx):

No problem